A balloon is filled to a volume of 23.7L at a temperature of 27.1ºC. If the pressure in the balloon is measured to be 2.20 atm, how many moles of gas are contained inside the balloon?


1
Expert's answer
2021-03-03T01:58:53-0500

Volume, V = 23.7 L.

Temperature, T = 27.1 ℃ = (27.1 + 273) K

= 300.1 K.

Pressure, P = 2.20 atm.

From ideal gas law, the number of moles of gas, n = (P × V)/(R × T)

= (2.20 atm × 23.7 L )/(0.0821 L.amt.mol-1.K-1 × 300.1 K) = 2.12 moles.